evaluate the following integral
∫ 1 + tan x / x + log sec x dx

Grade:12

1 Answers

Sanjeev Kumar
9 years ago
  • Let x​ + log Sec x = t
1 + 1/sec x ​×sec x tan x dx = dt
= 1+ tan x dx =dt
equating the above value in
∫1 + tan x / x + log sec x dx
we get
∫dt/t
= log t
= log (x + sec x)